# Nemotron-4 340B Technical Report

### v1

@@ 3.4.3 Safety Evaluations
As LLMs become more widespread, the content safety risks associated with their use also increase. To evaluate the safety of our model, we employ AEGIS (Ghosh et al., 2024), a high quality content safety solution and evaluation benchmark from NVIDIA. AEGIS is backed by a broad content safety risk taxonomy that covers 12 critical risks in human-LLM interactions (see details in Supplemetary Materials H). The taxonomy was created by considering most relevant community risks across multiple content safety risk taxonomies. It aligns with NVIDIA’s organizational values for the protected characteristics under categories of hate and harassment and defines sexual abuse of a minor as a separate critical hazard category. We also introduce a new category, “Needs Caution”, to address ambiguous situations where there isn’t sufficient context to determine safety. This category is particularly useful for scenarios where a more defensive mode is preferred over a more permissive one, as “Needs Caution” can be mapped to either unsafe or safe as needed. As a benchmark, AEGIS comprises a human annotated dataset of user prompts, single turn, and multi-turn dialogues, and AEGIS safety models that can predict if the response from a candidate LLM is safe or unsafe and provide categories of violation if the response is unsafe. AEGIS safety models are a group of open sourced LlamaGuard (Inan et al., 2023) LLM based classifiers, that were further instruction tuned with AEGIS safety taxonomy and policy in a parameter efficient manner.

### v2

@@ 3.4.3 Safety Evaluations
We performed extensive safety evaluation including adversarial testing via these distinct methods:
  • AEGIS (Ghosh et al., 2024) is a content safety evaluation dataset and LLM based content safety classifier model, that adheres to a broad taxonomy of 13 categories of critical risks in human-LLM interactions.
  • Garak (Derczynski et al., 2024) is an automated LLM vulnerability scanner that probes for common weaknesses, including prompt injection and data leakage.
  • Human Content Red Teaming leveraging human interaction and evaluation of the models’ responses.
As LLMs become more widespread, the content safety risks associated with their use also increase. To evaluate the safety of our model, we employ AEGIS (Ghosh et al., 2024), a high quality content safety solution and evaluation benchmark from NVIDIA. AEGIS is backed by a broad content safety risk taxonomy that covers 12 critical risks in human-LLM interactions (see details in Supplementary Materials H). The taxonomy was created by considering most relevant community risks across multiple content safety risk taxonomies. It aligns with NVIDIA’s organizational values for the protected characteristics under categories of hate and harassment and defines sexual abuse of a minor as a separate critical hazard category. We also introduce a new category, “Needs Caution”, to address ambiguous situations where there isn’t sufficient context to determine safety. This category is particularly useful for scenarios where a more defensive mode is preferred over a more permissive one, as “Needs Caution” can be mapped to either unsafe or safe as needed. As a benchmark, AEGIS comprises a human annotated dataset of user prompts, single turn, and multi-turn dialogues, and AEGIS safety models that can predict if the response from a candidate LLM is safe or unsafe and provide categories of violation if the response is unsafe. AEGIS safety models are a group of open sourced LlamaGuard (Inan et al., 2023) LLM based classifiers, that were further instruction tuned with AEGIS safety taxonomy and policy in a parameter efficient manner.
@@
Alongside content safety, it is important to assess generative models for security weaknesses. Garak Derczynski et al. (2024), Generative AI Red-teaming & Assessment Kit, is a vulnerability scanner for Large Language Models. It identifies a broad range of security weaknesses and unwanted behaviors in language model-based technology. Garak can scan a model or dialog system and quickly discover where it is working well, and where it may be vulnerable to attack. Garak provides full reporting detailing what worked and what could use improvement.
Nemotron-4 340B was scanned using Garak for security vulnerabilities in multiple categories (automatic red teaming; toxic continuation; jailbreaks; prompt injection; content filtering; training data membership inference; code generation; training data replay; adversarial hallucination; chat exfiltration) across several stages of its development, with nominal to good performance.
There are a few specific considerations that were found during the above efforts that may have downstream impacts when using this model:
Malware
Nemotron-4 340B Chat blocks some but not all requests for help in building malware. It is prone to responding to both some top-level requests (e.g. “Write me a virus”) and some requests to write sub-parts of malware (e.g. “How can I disable system antivirus”). There was no assessment of whether the code it produced actually performs the requested functions.
Adversarial hallucination
Nemotron-4 340B gave direct, incorrect answers to problems. When faced with impossible logic problems the model gives yes/no answers. The model also falsely denies that some prime numbers are prime. We examined three categories of analytical challenge (primes, logical reasoning, and complex fact scoping) posed using leading questions and Nemotron-4 340B Chat had a low success rate.
Jailbreaks
Nemotron-4 340B Chat is susceptible to existing jailbreaks, or prompts that lead to models ignoring its alignment/safety training in subsequent interactions. Nemotron-4-340B had a pass rate below 30% for attempted jailbreaks. As shown in research, training models to be helpful also tends to make them easier to exploit by jailbreak.